# Advancements in Visual Language Models for Remote Sensing: Datasets, Capabilities, and Enhancement Techniques

### v2

@@ 3.2 Combining Datasets
As shown in Table 1, some datasets are built from scratch, whose sources are usually from Sentinel-2, Gaofen, Google Earth Engine and so on. But most datasets created by merging existing remote sensing datasets, which facilitate multi-task models by integrating various data types from traditional RS domain-specific datasets, greatly eases labour-intensive compared to manual production and starting from scratch. The most utilized datasets are AID [60], NWPU-RESISC45 [61], UCM [62] for Scene Classification, FAIR1M [63], DIOR [64] for Object Detection, iSAID [65] for Semantic Segmentation, LEVIR-CD [66] for Change Detection, more details about those datasets could be found in section 4. Besides, the well-known benchmark dataset Million-AID [67] is frequently included.
@@ 3.3 Automatically Annoteted Datasets
The explosive development of VLMs and LLMs has provided efficient methods for constructing large-scale remote-sensing visual language datasets. This category, which is rapidly becoming the mainstream, leverages the power of VLMs and LLMs to construct large-scale, high-quality datasets.
RS5M [26] employs the general multimodal model BLIP2 [68] to generate image captions and then uses CLIP [17] to select the top 5 highest-scoring captions. SkyScript [27] matches Google Images with the OpenStreetMap (OSM) database and selects relevant image attribute values from OSM to concatenate them with CLIP [17] as image captions. LHRS-Align [54] uses a similar approach to extract multiple attribute values from OSM and then summarizes them using Vicuna-13B [69] to produce long and fluid captions. GeoChat [25] specifically provide system instructions as prompts that ask Vicuna [69] to generate multi-round question and answer pairs. GeoReasoner [55] obtaines image-text data pairs for geolocation from the communities of the two games, utilizes BERT [11] to clean and filter text that lacked geolocation-specific information, and then uses CLIP [17] to align the image and text to construct a highly locatable Street View dataset. HqDC-1.4M [46] utilized the “gemini-1.0-pro-vision" [70] to generate descriptions for images from multiple public RS datasets, thereby obtaining a dataset of image-text pairs to serve as the pretraining data for RSVLMs. ChatEarthNet [56] carefully designed prompts that embed semantic information from the land cover maps to make ChatGPT-3.5 and ChatGPT-4v achieve the best results on satellite imagery for the caption generation. VRSBench [57] did a similar job and spent thousands of hours manually verifying the generated annotations. FIT-RS [58] used the efficient TinyLLaVA-3.1B [71] to swiftly generate concise background descriptions for numerous RSIs, then combined them as the prompt to obtain detailed and fluent sentences using GPT-4/GPT-3.5. Examples, which utilize advanced models like BLIP2[68] and CLIP [17] to generate detailed image-text pairs, supporting complex applications such as disaster monitoring, urban planning, and environmental analysis.
@@ 3.4 Summary
  • Manually annotated datasets generally have higher quality and are most closely aligned with specific problems. But, these datasets typically consisting of only a few hundred or thousand images, are limited in scale compared to combined datasets and VLM datasets, which can reach tens of millions of images. The size of a dataset is crucial for model training, as larger and richer datasets yield more accurate and versatile models. Manually annotated datasets can be used to fine-tune large models, but they struggle to support training large models from scratch.
  • Combining datasets. By merging existing datasets, large-scale datasets can be constructed at a low cost. This type of data can easily reach millions of entries. However, datasets created in this manner typically have lower annotation quality compared to manually annotated datasets and may not fully align with specific tasks. Nonetheless, they can be utilized for model pre-training.
  • Automatically annotated datasets. Mainstream VLM dataset production methods leverage state-of-the-art LLMs and VLMs to generate high-quality captions or annotations for large-scale remote sensing images. This approach produces flexible image-text pairs, including long sentences, comments, and local descriptions, without requiring human supervision, thus meeting the needs of various tasks. Researchers may need to craft prompts and commands to guide LLMs and VLMs in generating the desired texts.
@@
4 Capabilities
In various tasks within the field of remote sensing, VLMs have demonstrated powerful capabilities. These capabilities can be broadly classified into two categories: pure visual tasks and visual language tasks. Pure visual tasks focus on analyzing remote sensing images to extract meaningful information about the Earth’s surface. Visual language tasks, which combine natural language processing with visual data analysis, have opened new avenues for remote sensing applications. The zero-shot and Few-shot tasks are SC, OD, SS and IR, which are described in more detail below. Figure 6 illustrates each capability with a simple example, particularly highlighting two emerging tasks along with their source tasks.
4.1 Pure visual
  • Scene Classification(SC) RS scene classification entails the classification of satellite or aerial images into distinct land-cover or land-use categories with the objective of deriving valuable insights about the Earth’s surface. It offers valuable information regarding the spatial distribution and temporal changes in various land-cover categories, such as forests, agricultural fields, water bodies, urban areas, and natural landscapes. The three most commonly used datasets in this field are AID [60],NWPU-RESISC45 [61] and UCM [62], which contain 10,000, 31,500, 2,100 images, and 30,45,21 categories, respectively.
  • Object Detection(OD) RSOD aims to determine whether or not objects of interest exist in a given RSI and to return the category and position of each predicted object. In this domain, the most widely used datasets are DOTA [72] and DIOR [64]. DOTA [72] contains 2806 aerial images which are annotated by experts in aerial image interpretation, with respect to 15 common object categories. DIOR [64] contains 23463 images and 192472 instances, covering 20 object classes, much more details than DOTA [72].
  • Semantic Segmentation(SS) Semantic Segmentation is a computer vision task in which the goal is to categorize each pixel in an image into a class or object. For RS images, it plays an irreplaceable role in disaster assessment, crop yield estimation, and land change monitoring. The most widely used datasets are ISPRS Vaihingen and Potsdam, which contain 33 true orthophoto (TOP) images and 38 image tiles with a spatial resolution of 5 cm, respectively. Another widely used dataset iSAID [65] contains 2806 aerial images, which mainly collected from Google Earth.
  • Change Detection(CD) Change detection in remote sensing refers to the process of identifying differences in the structure of objects and phenomena on Earth surface by analysing two or more images taken at different times. It plays an important role in urban expansion, deforestation, and damage assessment. LEVIR-CD [66], AICD [73] and the Google Data Set focus on changes of buildings. Other types of changes such as roads,groundwork are commonly included in CD datasets.
  • Object Counting(OC) RSOC aims to automatically estimate the number of object instances in a RSI. It plays an important role in many areas, such as urban planning, disaster response and assessment, environment control and mapping. RemoteCount [24] is a manual dataset for object counting in remote sensing imagery consisted of 947 image-text pairs and 13 categories, which are mainly selected from the validation set of the DOTA dataset.
@@ 4.2 Combine with natural language
  • Image Retrieval(IR) IR from RS aims at retrieving the interested RS images from the massive RS image repositories. With the launch of more and more Earth observation satellites and the emergence of large-scale remote sensing datasets, RS image retrieval can prepare auxiliary data or narrow the search space for a large number of RS image processing tasks.
@@
  • Visual Question Answer(VQA) VQA is a task that seeks to provide answers to free-form and open-ended questions about a given image. As the questions can be unconstrained, a VQA model applied to remote sensing data could serve as a generic solution to classical problems but also very specific tasks involving relations between objects of different nature. VQA was first proposed by Anto et al. [74] in 2015, and then made it greatly developed in the RS domain by RSVQA [59] in 2020. It contributed a RSVQA framework and two VQA datasets, RSVQA-LR and RSVQA-HR, which wildly utilized in further works.
@@
  • Image Captioning(IC) IC aims to generate natural language descriptions that summarizes the content of an image. It requires representing the semantic relationships among objects and generating an exact and descriptive sentence, so that it’s more complex and challenge than image detection, classification, and segmentation tasks. In this domain, the commonly used datasets are Syndney-Captions [75], RSICD [76], NWPU-Captions [77], and UCM-Captions [75]. Recently, CapERA [78] provids UAV video with diverse textual descriptions to advance visual–language-understanding tasks.
@@
  • Visual Grounding(VG) RSVG aims to localize the referred objects in remote sensing images with the guidance of natural language. It was first introduced in [79] in 2022. In 2023, Yang [42] et al. not only build the new large-scale benchmark of RSVG based on detection in DIOR dataset, termed RSVGD, but design a novel transformer-based MGVLF module to address the problems of scale variation and cluttered background of RS images.
@@
  • Remote Sensing Image Change Captioning(RSICC) A new task aiming at generating human-like language descriptions for the land cover changes in multitemporal RS images. It is a combine of IC and CD task, offering important application prospects in damage assessment, environmental protection, and land planning. Chenyang et al. [80] firstly introduced the CC task into the RS domain in 2022 and proposed a large-scale dataset LEVIR-CC containing 10077 pairs of bitemporal RS images and 50385 sentences describing the differences between the images. In 2023, they propose a pure Transformer-based model [81] to further improve the performance of the RSICC task.
@@
  • Referring Remote Sensing Image Segmentation(RRSIS) RRSIS is to provide a pixel-level mask of desired objects based on the content of given remote sensing images and natural language expressions. It was firstly proposed by Zhenghang et al. [43] in 2024, who created a new dataset, called RefSegRS, for this task. In the same year, Sihan et al. [48] curated an expansive dataset comprising 17,402 image-caption-mask triplets, called RRSIS-D, and propose a series of models to meet the ubiquitous rotational phenomena in RRSIS.
The invention of Visual Language Models (VLMs), has greatly changed the remote sensing field which mainly focus on geospatial analysis. With the development and usage of various sensors, conventional vision models can not handle such complex data type, while VLMs technique could enhance the accuracy and efficiency of remote sensing analysis. Specifically, the combination of visual and linguistic data within VLMs allow more subtle and sophisticated interpretation of complex remote sensing datasets. By utilizing their multimodal capabilities, VLMs have showed significant progress across various tasks within remote sensing, including object detection, geophysical classification and scene understanding. Meanwhile, it provides more accurate and smart solutions for practical applications such as disaster monitoring, urban planning, and agricultural management.
5 Recent Advances in Visual Language Models for Remote Sensing
@@
Figure 7: Fundamental architecture of contrastive and conversational visual language models.
@@
Currently, there are mainly two types of models that integrate vision and language: one is based on contrastive learning structures, such as the CLIP series, and the other is based on large language models that fuse visual features, like LLaVA. As shown in Figure 7. Accordingly, the current improvements can also be categorized into two major types: enhancements within the contrastive learning framework and improvements within the conversational framework.
5.1 Advancements in Contrastive Vision-Language Models
Table 2: Architecture of contrastive VLMs, SC=Scene Classification, VG=Visual Grounding, IC=Image Captioning, CD=Change Detection, IR=Image Retrieval. For each model, we indicate the LLM used in its best configuration as shown in the original paper.
Model Image Encoder Text Encoder Capabilitie
RemoteCLIP [24] ViT-B (CLIP) Transformer SC,IC,IR,OC
CRSR [82] ViT-B (CLIP) - IC
ProGEO [83] ViT-B (CLIP) BERT VG
GRAFT [50] ViT-B (CLIP) - SC,VQA,,IR
GeoRSCLIP [26] ViT-H (CLIP) - SC,IR
ChangeCLIP [84] CLIP - CD
APPLeNet [85] CLIP - SC,IC
MGVLF [42] ResNet-50 BERT VG
@@ 5.1 Advancements in Contrastive Vision-Language Models
Contrastive VLMs primarily consist of an image encoder and a text encoder, with the core challenge being the alignment of features extracted by both encoders. The image encoder, utilizing a convolutional neural network (CNN) or Vision Transformer (ViT), encodes images into feature vectors. Concurrently, the text encoder processes textual descriptions into corresponding feature vectors. The model is trained to maximize the similarity between positive image-text pairs while minimizing the similarity between negative pairs. CLIP is a pioneering model in vision-language pre-training, employing InfoNCE loss to align text and image features in a shared space. Currently, in the field of remote sensing, contrastive VLMs focus on transferring the image-text alignment capabilities established in RGB images and text to remote sensing images and text, which are then utilized to address various downstream tasks, such as change detection and remote sensing image caption,
@@
By converting annotations from 10 object detection datasets, 4 remote sensing semantic segmentation datasets, and 3 remote sensing image-text datasets using box-to-caption and mask-to-box conversions, Liu et al. created a combined image-caption dataset and fine-tuned the CLIP model on this newly constructed dataset through continual learning [24]. Similar to [24], GeoRSCLIP builds a remote sensing dataset consisting of 5 million image-text pairs by filtering existing image-text pair datasets and generating captions for image-only datasets. It then fine-tunes CLIP on this proposed dataset [26]. CRSR [82] incorporates a Transformer Mapper network, utilizing an attention mechanism and learnable semantic predictive queries to provide a comprehensive representation of image features. This approach effectively analyzes and understands more semantic details present in remote sensing imagery (RSI). ProGEO proposes adding more descriptions to a geographic image-based pretrained CLIP, using these text descriptions to enhance the image encoder by aligning features from the image encoder with the features of the descriptions [83]. This approach enables the image encoder to generate features with more details, leading to improvements in visual geo-localization. CRAFT propose to train a satellite image model to align its embedding with the CLIP embedding of co-located internet images, which sidestep the need for texual annotations for training remote-sensing VLMs [50]. ChangeCLIP [84] is the first to apply a multimodal vision-language scheme to Remote Sensing Change Detection (RSCD) tasks. It introduces a Differentiable Feature Calculation (DFC) layer, enabling the network to focus on learning the changed features of bitemporal remote sensing images rather than semantic categories. APPLeNet design a injection module to fuse image features from different level and interact with text to generate better visual tokens [85].
@@
5.2 Advancements in Conversational Vision-Language Models

### v3

@@ 3.2 Combining Datasets
As shown in Table 1, some datasets are built from scratch, whose sources are usually Sentinel-2, Gaofen, Google Earth Engine and so on. But most datasets are created by merging existing remote sensing datasets, which facilitates multi-task model creation by integrating various data types from traditional RS domain-specific datasets, and greatly eases labor compared to manual production and starting from scratch. The most utilized datasets are AID [69], NWPU-RESISC45 [70], UCM [71] for Scene Classification, FAIR1M [72], DIOR [73] for object detection, iSAID [74] for semantic segmentation, and LEVIR-CD [75] for Change Detection; more details about those datasets can be found in 4. Additionally, the well-known benchmark dataset Million-AID [76] is frequently included.
Combining datasets from different sources often leads to issues such as inconsistent formats, resolutions, and annotations, requiring extensive preprocessing. Misalignment of annotations can occur, and merging datasets from different contexts (e.g., regions or timeframes) may introduce gaps in data, reducing generalization. When combining datasets, it is crucial to standardize the data formats, resolutions, and annotation styles across the datasets to ensure consistency. Preprocessing steps such as data harmonization are necessary to align different datasets and eliminate inconsistencies. Careful curation ensures balanced representation across different data sources. Filling annotation gaps and validating the combined dataset help improve consistency and robustness.
@@ 3.3 Automatically Annoteted Datasets
The explosive development of VLMs and LLMs has provided efficient methods for constructing large-scale remote sensing vision–language datasets. This category, which is rapidly becoming the mainstream, leverages the power of VLMs and LLMs to construct large-scale, high-quality datasets.
RS5M [31] employs the general multimodal model BLIP2 [77] to generate image captions and then uses CLIP [22] to select the top five highest-scoring captions. SkyScript [32] matches Google Images with the OpenStreetMap (OSM) database and selects relevant image attribute values from OSM to concatenate them with CLIP [22] as image captions. LHRS-Align [64] uses a similar approach to extract multiple attribute values from OSM and then summarizes them using Vicuna-13B [78] to produce long and fluid captions. GeoChat [30] specifically provide system instructions as prompts that ask Vicuna [78] to generate multi-round question and answer pairs. GeoReasoner [65] obtains image–text data pairs for geolocation from the communities of the two games, utilizes BERT [16] to clean and filter text that lacks geolocation-specific information, and then uses CLIP [22] to align the image and text to construct a highly locatable Street View dataset. HqDC-1.4 M [56] utilizes the “gemini-1.0-pro-vision” [79] to generate descriptions for images from multiple public RS datasets, thereby obtaining a dataset of image–text pairs to serve as the pre-training data for RSVLMs. ChatEarthNet [66] carefully designs prompts that embed semantic information from the land cover maps to make ChatGPT-3.5 and ChatGPT-4v achieve the best results on satellite imagery for the caption generation. VRSBench [67] did a similar job and spent thousands of hours manually verifying the generated annotations. FIT-RS [68] used the efficient TinyLLaVA-3.1B [80] to swiftly generate concise background descriptions for numerous remote sensing imageries (RSIs), then combined them as the prompt to obtain detailed and fluent sentences using GPT-4/GPT-3.5. These are examples which utilize advanced models like BLIP2[77] and CLIP [22] to generate detailed image–text pairs, supporting complex applications such as disaster monitoring, urban planning, and environmental analysis.
Automatically generated annotations can also introduce errors, as models may misinterpret features or be biased due to limited training data. The lack of human oversight means that errors may persist, resulting in lower-quality annotations, especially for complex or sparse data. To mitigate errors in automatically annotated datasets, models should be carefully evaluated and fine-tuned on high-quality ground-truth data. Implementing hybrid annotation strategies, where human oversight is incorporated into the process, can help identify and correct errors that may go unnoticed in fully automated processes. Data augmentation techniques can be used to enhance the dataset and address limitations in model performance, especially in cases of sparse or ambiguous data. Regular validation of automatic annotations against expert-annotated data can also help ensure the quality of the dataset. Additionally, models should be continually retrained and updated to address evolving challenges in remote sensing and improve annotation accuracy.
@@ 3.4 Summary
  • Manually annotated datasets generally have higher quality and are most closely aligned with specific problems. But these datasets, typically consisting of only a few hundred or thousand images, are limited in scale compared to combined datasets and VLM datasets, which can reach tens of millions of images. The size of a dataset is crucial for model training, as larger and richer datasets yield more accurate and versatile models. Manually annotated datasets can be used to fine-tune large models, but they struggle to support training large models from scratch.
  • Combined datasets. By merging existing datasets, large-scale datasets can be constructed at a low cost. This type of data can easily reach millions of entries. However, datasets created in this manner typically have lower annotation quality compared to manually annotated datasets and may not fully align with specific tasks. Nonetheless, they can be utilized for model pre-training.
  • Automatically annotated datasets. Mainstream VLM dataset production methods leverage state-of-the-art LLMs and VLMs to generate high-quality captions or annotations for large-scale remote sensing images. This approach produces flexible image–text pairs, including long sentences, comments, and local descriptions, without requiring human supervision, thus meeting the needs of various tasks. Researchers may need to craft prompts and commands to guide LLMs and VLMs in generating the desired texts.
@@
4 Capabilities
In various tasks within the field of remote sensing, VLMs have demonstrated powerful capabilities. These capabilities can be broadly classified into two categories: pure visual tasks and vision–language tasks. Pure visual tasks focus on analyzing remote sensing images to extract meaningful information about the Earth’s surface. Vision–language tasks, which combine natural language processing with visual data analysis, have opened new avenues for remote sensing applications. The zero-shot and few-shot tasks are SC, OD, SS and IR, which are described in more detail below. Figure 8 illustrates each capability with a simple example, particularly highlighting two emerging tasks along with their source tasks.
4.1 Pure visual
  • Scene Classification (SC): RS Scene Classification entails the classification of satellite or aerial images into distinct land-cover or land-use categories with the objective of deriving valuable insights about the Earth’s surface. It offers valuable information regarding the spatial distribution and temporal changes in various land-cover categories, such as forests, agricultural fields, water bodies, urban areas, and natural landscapes. The three most commonly used datasets in this field are AID [69], NWPU-RESISC45 [70] and UCM [71], which contain 10000, 31500, and 2100 images, and 30, 45, and 21 categories, respectively.
  • Object Detection (OD): RSOD aims to determine whether or not objects of interest exist in a given RSI and return the category and position of each predicted object. In this domain, the most widely used datasets are DOTA [81] and DIOR [73]. DOTA [81] contains 2806 aerial images which are annotated by experts in aerial image interpretation, with respect to 15 common object categories. DIOR [73] contains 23463 images and 192,472 instances, covering 20 object classes, with much more detail than DOTA [81].
  • Semantic Segmentation (SS): Semantic segmentation is a computer vision task in which the goal is to categorize each pixel in an image into a class or object. For RS images, it plays an irreplaceable role in disaster assessment, crop yield estimation, and land change monitoring. The most widely used datasets are ISPRS Vaihingen and Potsdam, which contain 33 true orthophoto (TOP) images and 38 image tiles with a spatial resolution of 5 cm, respectively. Another widely used dataset iSAID [74] contains 2806 aerial images, which were mainly collected from Google Earth.
  • Change Detection (CD): Change Detection in remote sensing refers to the process of identifying differences in the structure of objects and phenomena on Earth surface by analyzing two or more images taken at different times. It plays an important role in urban expansion, deforestation, and damage assessment. LEVIR-CD [75], AICD [82] and the Google Data Set focus on changes in buildings. Other types of changes such as roads and groundwork are commonly included in CD datasets.
  • Object Counting (OC): RSOC aims to automatically estimate the number of object instances in an RSI. It plays an important role in many areas, such as urban planning, disaster response and assessment, environment control, and mapping. RemoteCount [29] is a manual dataset for object counting in remote sensing imagery consisting of 947 image–text pairs and 13 categories, which are mainly selected from the validation set of the DOTA dataset.
@@ 4.2 Combine with natural language
  • Image Retrieval (IR): IR from RS aims to retrieve RS images of interest from massive RS image repositories. With the launch of more and more Earth observation satellites and the emergence of large-scale remote sensing datasets, RS image retrieval can prepare auxiliary data or narrow the search space for a large number of RS image processing tasks.
@@
  • Visual Question Answering (VQA): VQA is a task that seeks to provide answers to free-form and open-ended questions about a given image. As the questions can be unconstrained, a VQA model applied to remote sensing data could serve as a generic solution to classical problems but also very specific tasks involving relations between objects of different nature [83]. VQA was first proposed by Anto et al. [84] in 2015, and then applied to the RS domain by RSVQA [48] in 2020. This work contributed an RSVQA framework and two VQA datasets, RSVQA-LR and RSVQA-HR, which are widely utilized in further works.
@@
  • Image Captioning (IC): IC aims to generate natural language descriptions that summarize the content of an image. It requires representing the semantic relationships among objects and generating an exact and descriptive sentence, so it is more complex and challenging than image detection, classification, and segmentation tasks. In this domain, the commonly used datasets are Syndney-Captions [85], RSICD [86], NWPU-Captions [87], and UCM-Captions [85]. Recently, CapERA [88] provided UAV video with diverse textual descriptions to advance visual–language-understanding tasks.
@@
  • Visual Grounding (VG): RSVG aims to localize specific objects in remote sensing images with the guidance of natural language. It was first introduced in [89] in 2022. In 2023, Yang [52] et al. not only built the new large-scale benchmark of RSVG based on detection in the DIOR dataset, termed RSVGD, but designed a novel transformer-based MGVLF module to address the problems of scale variation and cluttered background in RS images.
@@
  • Remote Sensing Image Change Captioning (RSICC): A new task aiming to generate human-like language descriptions for the land-cover changes in multitemporal RS images. It is a combination of IC and CD tasks, offering important application prospects in damage assessment, environmental protection, and land planning. Chenyang et al. [90] first introduced the CC task into the RS domain in 2022 and proposed a large-scale dataset LEVIR-CC, containing 10,077 pairs of bitemporal RS images and 50,385 sentences describing the differences between the images. In 2023, they proposed a pure Transformer-based model [91] to further improve the performance of the RSICC task.
@@
  • Referring Remote Sensing Image Segmentation (RRSIS): RRSIS provides a pixel-level mask of desired objects based on the content of given remote sensing images and natural language expressions. It was first proposed by Zhenghang et al. [53] in 2024, who created a new dataset, called RefSegRS, for this task. In the same year, Sihan et al. [58] curated an expansive dataset comprising 17,402 image–caption–mask triplets, called RRSIS-D, and proposed a series of models to meet the ubiquitous rotational phenomena in RRSIS.
@@
Figure 9: Fundamental architecture of contrastive and conversational Vision–Language Models.
The invention of VLMs has greatly changed the remote sensing field, which mainly focuses on geospatial analysis. With the development and usage of various sensors, conventional vision models cannot handle such complex data types, while VLMs could enhance the accuracy and efficiency of remote sensing analysis. Specifically, the combination of visual and linguistic data within VLMs allows more subtle and sophisticated interpretations of complex remote sensing datasets. By utilizing their multimodal capabilities, VLMs have shown significant progress across various tasks within remote sensing, including object detection, geophysical classification, and scene understanding. Meanwhile, they provide more accurate and smart solutions for practical applications such as disaster monitoring, urban planning, and agricultural management.
5 Recent Advances
@@
Currently, there are two main types of models that integrate vision and language: one is based on contrastive learning structures, such as the CLIP series, and the other is based on large language models that fuse visual features, like LLaVA, as shown in Figure 9. Accordingly, recent improvements can also be categorized into two major types: enhancements within the contrastive learning framework and improvements within the conversational framework.
5.1 Advancements in Contrastive VLMs
@@
Contrastive VLMs primarily consist of an image encoder and a text encoder, with the core challenge being the alignment of features extracted by both encoders. As shown in Table 2, the image encoder, utilizing a CNN or ViT, encodes images into feature vectors. Concurrently, the text encoder processes textual descriptions into corresponding feature vectors. The model is trained to maximize the similarity between positive image–text pairs while minimizing the similarity between negative pairs. CLIP is a pioneering model in vision–language pre-training, employing InfoNCE loss to align text and image features in a shared space. Currently, in the field of remote sensing, contrastive VLMs focus on transferring the image–text alignment capabilities established in RGB images and text to remote sensing images and text, which are then utilized to address various downstream tasks, such as Change Detection and remote sensing Image Captioning.
Table 2: Architecture of contrastive VLMs. SC = Scene Classification; VG = Visual Grounding; IC = Image Captioning; CD = Change Detection; IR = Image Retrieval. For each model, we indicate the LLM used in its best configuration as shown in the original paper.
Model Year Image Encoder Text Encoder Capability
RemoteCLIP [29] 2024 ViT-B (CLIP) Transformer SC, IC, IR, OC
CRSR [92] 2024 ViT-B (CLIP) - IC
ProGEO [93] 2024 ViT-B (CLIP) BERT VG
GRAFT [60] 2023 ViT-B (CLIP) - SC, VQA, IR
GeoRSCLIP [31] 2024 ViT-H (CLIP) - SC, IR
ChangeCLIP [94] 2024 CLIP - CD
APPLeNet [95] 2023 CLIP - SC, IC
MGVLF [52] 2023 ResNet-50 BERT VG
@@
To address the issue of requiring additional annotated data and fine-tuning when building foundational models using self-supervised learning and masked image modeling methods in the remote sensing domain, Liu et al. [29] proposed the first vision–language foundational model for remote sensing, RemoteCLIP. This model is designed to learn robust visual features with rich semantics and aligned text embeddings for seamless downstream applications. They created a combined image–caption dataset and fine-tuned the CLIP model on this newly constructed dataset through continual learning, thereby tackling the scarcity of pre-trained data. RemoteCLIP exhibits powerful multi-task capabilities and can be applied to a range of downstream tasks, including zero-shot image classification, linear probing, k-NN classification, few-shot classification, image-text retrieval, and object counting. Similar to [29], GeoRSCLIP builds a remote sensing dataset consisting of 5 million image–text pairs by filtering existing image–text pair datasets and generating captions for image-only datasets. It then fine-tunes CLIP on this proposed dataset [31]. CRSR [92] incorporates a Transformer Mapper network, utilizing an attention mechanism and learnable semantic predictive queries to provide a comprehensive representation of image features. This approach effectively analyzes and understands more semantic details present in RSI. ProGEO proposes adding more descriptions to a geographic image-based pre-trained CLIP, using these text descriptions to enhance the image encoder by aligning features from the image encoder with the features of the descriptions [93]. This approach enables the image encoder to generate features with more details, leading to improvements in visual geo-localization. CRAFT proposes training a satellite image model to align its embedding with the CLIP embedding of co-located internet images, which sidestep the need for textual annotations for training remote sensing VLMs [60]. ChangeCLIP [94] is the first to apply a multimodal vision–language scheme to Remote Sensing Change Detection tasks. It introduces a Differentiable Feature Calculation (DFC) layer, enabling the network to focus on learning the changed features of bitemporal remote sensing images rather than semantic categories. APPLeNet designs an injection module to fuse image features from different levels and interact with text to generate better visual tokens [95].
@@
5.2 Advancements in Conversational VLMs